# A Comprehensive Survey on Data Augmentation

### v3

@@ VI-A Vanilla Generation
A straightforward way to produce augmented data from the entire dataset is to learn the data distribution of existing datasets with neural networks and generate new samples that fall in that distribution. Some typical solutions are Autoencoder (AE), Generative Adversarial Networks (GAN), Large Language Models (LLMs), and Generative Diffusion Models.
@@ VI-A1 Autoencoder
An Autoencoder (AE) learns from the dataset and extracts the essential structure, which can be used to generate a more robust dataset. It is widely used for tabular data augmentation. [170] tries out four types of autoencoder, concluding that all of these autoencoders can generate data sets that produce better task performance, while the variational autoencoder (VAE) provides the most robust result. Some other works [171, 172] also use VAE as their data generation approach. SDAT [171] adds extra noise to the latent space of its VAE to obtain augmented samples. These samples, along with the original ones, are then trained under a semi-supervised framework. [172] combines VAE with active learning. The samples are iteratively selected by active learning, labeled by an oracle, reconstructed by the VAE, and discriminated and classified.
@@ VI-A2 Generative Adversarial Networks
The rise of Generative Adversarial Networks (GANs) [173] makes it possible for humans to generate new images on a large scale using computers. This nature also leads to a new path of data augmentation: creating new data samples instead of modifying or combining old ones. It can produce considerably more augmented data and is especially useful for oversampling and alleviating the problem of class imbalance.
For image data, GANs have now been widely used as a data augmentation technique [12], especially in the medical field [174, 175], where data is to some extent scarce. Apart from simply enlarging the dataset, GAN can be used in other ways such as removing background noise and cleaning the dataset [176, 177, 178]. It can also generate different views or poses of the object based on the original image [179], and thus add more variation to the dataset.
Tabular data augmentation is another case where GANs play an important role. Table-GAN [180] mainly considers the privacy issue, and experiments show that the augmented table given by GAN is compatible with downstream machine learning models and produces similar performance. CTGAN [181] focuses on the class imbalance problem and proposes a conditional generator that accounts for the imbalance in categorical columns. [182] shares a similar goal and method with CTGAN [181] but employs a more sophisticated loss function that better incorporates the discriminator and the auxiliary classifier. ITS-GAN [183] proposes a framework that aims to preserve functional dependencies in tables. It models functional dependencies with autoencoders and includes a term to characterize the set of functional dependency constraints in the generator loss function. Arguing that GANs are difficult to train on small datasets, ConvGeN [184] designed a new mechanism that requires the GAN’s generator and discriminator to cooperate, that the generator generates batches of a convex combination of minority neighborhood samples and the discriminator classifies these augmented minority samples against majority batches.
GANs are also widely used to generate virtual time-series data. [185] uses an RGAN and an RCGAN to produce realistic medical time-series data. TimeGAN [186] combines supervised and unsupervised losses to train its autoencoding components and adversarial components. [187] propose WaveGAN and SpecGAN to generate time-series data. WaveGAN is used on the waveform, while SpecGAN is used on the spectrogram. SpecGAN demonstrates how GAN is applied to generate time-series data on the frequency domain.
@@ VI-A3 Large Language Models
Recent advances in large language models (LLM) and generative pre-training (GPT) [188] models have demonstrated unprecedented power in text understanding, classification, and generation tasks. Apart from supporting intelligent chat-bots that help people in their daily lives, the capacity of these state-of-the-art language generation models establishes a new paradigm of text data augmentation.
LAMBADA [189] is among the first attempts to use a state-of-the-art language generator to synthesize new text data for data augmentation, and its procedure is as follows. LAMBADA first trains a baseline classifier from a small dataset, and then it fine-tunes the GPT-2 [190] model. A set of labels is thereafter fed to the fine-tuned language model, which generates a large dataset of labeled sentences. These sentences are evaluated by the classifier trained prior according to its confidence score, and top-ranked sentences are retained to compose the synthesized dataset. Similarly, [191] integrates GPT-2 in its text data augmentation process and improves the performance of active learning on tiny datasets. It uses the language model to generate a tree of tokens as candidates and construct a sentence from them. The set of new sentences is repeatedly refined by active learning, and the outcome is an augmented text dataset.
Some other research also applies GPT-2 to effectively augment low-resource text datasets, such as [192] and [193]. Despite their success, both of them point out the need for more sophisticated methods to generate text data. [192] for example, argues that human-in-the-loop for relabeling might be rewarding. [193] adopts human-in-the-loop for seed selection and further discusses the potential benefits of injecting expert knowledge for specialized domains. Notably, Data Boost [194] proposes a reinforcement learning framework to guide the process of text data augmentation with GPT-2 and succeeds in generating data of high quality and high similarity to the original data.
The rise of GPT-3 [195] and later LLMs is now providing more possibilities for text generative augmentation. Their use and improvements are studied in the latest research such as [196] and [197].
@@ VI-A4 Generative Diffusion Models
Despite that AEs and GANs can all be used to generate graph data, they face many problems in practice. VAE models require things like a massive graph-matching process, and GANs on graphs easily fall into mode collapse [198]. So, graph data augmentation resorts to another recently proposed technique, namely, Generative Diffusion Models.
The denoising diffusion model [199] is a generative paradigm that has become popular in recent years. Though not yet widely applied as an approach to graph data augmentation, the Generative Diffusion Model boasts enormous potential in graph generation and can possibly be adopted as a promising data augmentation technique. To implement diffusion models on graph data, some properties of the graph data must be properly treated. For example, diffusion models often rely on a continuous Gaussian noise process, but the graph structures are discrete [200]. DiGress [201] designs a discrete denoising diffusion model for graph generation and achieves state-of-the-art performance on molecular and non-molecular datasets. The use of Generative Diffusion Models for data augmentation especially benefits studies on molecules, proteins, and materials, whose data are often in the form of graph instances.
@@ VI-B1 Introducing Exogenous Datasets
Tabular data populates the web and by retrieving consistent and related tables, it is possible to generate an augmented table that incorporates extended information. A key consideration of Table Retrieval Augmentation is to design appropriate ways to search for ideal tables. Santos [204] leverages external knowledge bases, YAGO 4, and focuses on designing a union search technique to match desired target tables. HYTREL [205] predicts table similarities by building hypergraphs on top of the tables and uses a hypergraph encoder to calculate the encoding. LEKA [206] uses an LLM to extract the table’s descriptive information, retrieve it in the Kaggle database, and harmonize tables from different sources.
@@ VI-B2 Leveraging Expert Knowledge
Computer Graphic Modeling is to leverage expert knowledge on how objects in the existing dataset look like and the power of computer software to generate new image samples [207, 208, 209]. These datasets have several advantages over ordinary image datasets. First, it requires little human effort to collect data. Second, by applying different aspects, poses, and lighting conditions, computer graphics modeling can generate images with high diversity. Third, these images are all fully labeled with accurate ground truth.
Relational Structure Construction can enhance the potential of tabular data, which is limited by its simple nature compared with data from other modalities. Due to the lack of spatial or relational relationships, many popular augmentation approaches and network configurations are not suitable for tabular datasets. A solution is to construct and introduce a relational structure based on the tabular dataset. GOGGLE [210] proposes a solution to this problem. It learns the relational structure and corresponding functional relationships from tabular datasets and uses this basis to generate new samples. In other words, it learns a graph from the features in the table and obtains augmented data from that graph. PET [211] depicts multiple tabular data instances as a hyper-graph and uses the labels to help construct hyper-edges. These methods are based on external assumptions and introduce extra structural information beyond the tabular dataset.
@@ VII Discussions
Up to now this survey focuses on wherefrom the information in the augmented samples is derived. And there are many other perspectives to discuss data augmentation and understand how it works.
@@ VII-A Augmentation Evaluation Metrics
Most research on data augmentation uses the models’ performance on downstream tasks to evaluate the effect of data augmentation [13]. However, this criterion depends on many extraneous factors and may fail to reflect the influence of data augmentation methods on the data. So, there is a need for a set of evaluation metrics that can assess specific considerations of data augmentation. Two of the most critical metrics are affinity and diversity [161, 162, 214, 215, 216, 217, 125].
Affinity is also referred to as consistency, correctness, faithfulness, quality, similarity, or validity. It reflects if the augmented sample is semantically consistent with the original sample and preserves the original label. Model-centric evaluation [215] defines affinity as the ratio of model $m$’s performance on augmented validation dataset $D^{\prime}_{val}$ and that on the original validation dataset $D_{val}$, given by
[…]
where $\text{ACC}(m,D)$ stands for model $m$’s performance on evaluation dataset $D$. Data-centric evaluation, on the other hand, defines affinity as how much the augmented dataset deviates from the original dataset in the feature space. For example, [214] proposes Fréchet BERT Distance (FBD) to evaluate the deviation of two textual datasets, defined as:
[…]
where $m$ is the mean vector and $C$ is the covariance matrix.
Diversity reflects the breadth of patterns in the dataset and exposing the model to a diverse training dataset can improve its robustness. [218] considers a dataset as an ellipsoid and defines its diversity as the isocontour radius of the ellipsoid:
[…]
where $H$ is the dimension of the feature vectors and $\sigma_{i}$ is the standard deviation along the $i$-th feature. [219] proposes to evaluate dataset diversity by calculating the dispersion:
[…]
where $d(x_{i},x_{j})=1-\frac{f(x_{i}){\cdot}f(x_{j})}{\|f(x_{i})\|\|f(x_{j})\|}$.
Some empirical studies including [219], [217], and [215] investigate how affinity and/or diversity influence the effectiveness of data augmentation, and their real-world experimental results demonstrate the model performance takes the most advantage of data augmentation that achieves both high affinity and diversity.
@@ VII-B1 On Selecting Pixels
For Pixel Erasing, selecting target pixels is a necessity because it is impossible to mask out all pixels in the image. If erasing is performed by randomly selecting an area in the image, it is called Image-aware Random Erasing (IRE) [10]. In some tasks, the location of the target in the image is known, so, random erasing can be restricted in the object’s bounding box, known as Object-aware random erasing (ORE) [10]. Further argued in [220], the erased area within the bounding box should be limited to a manageable size and a part of the less informative background should also be erased. [220] proposed Region-aware Random Erasing (RRE), which introduces ORE with upper bounds for width and height in the bounding boxes, and then randomly erases the background. The same consideration occurs when performing Image Cropping and Mixup, and some importance-based strategies are used to preserve important contents in the images. For Image Cropping (IC), [74] uses entropy as the importance score to select the most importance regions, and [75] uses eye-tracking technique to locate important content. For Image Mixup, Co-mixup [221] and Puzzle mix [222] both use saliency to decide the mixing strength of different regions, potentially leading to better mixing performances. A summary of how these methods select target pixels are presented in Table III.
TABLE III: Region selection strategy for Pixel Erasing, Image Cropping and Mixing methods.
Name Ref Augmentation Strategy
@@
Cutout [9]
Image-aware RE [10]
Object-aware RE [10]
Region-aware RE [220]
Hide-and-seek [38]
Grid-Mask [39]
Entropy-bases IC [74]
Gaze-based IC [75]
Co-mixup [221]
Puzzle mix [222]
@@ VII-B2 On Selecting Words
Most value-based text transformation methods randomly select the word to be replaced, added, and deleted, but some methods considers the role a word plays in the sentence when deciding whether to impose perturbation.
For example, TEXTFOOLER [223] ranks words by their importance and replaces the important words with their synonyms. Syntax-aware data augmentation [48] argues that most previous text data augmentation approaches are only computation-motivated and cannot assure the good quality of the generated text. Instead of simply modifying words with the same probability, Syntax-aware data augmentation calculates the probability of being modified for each word based on its syntactic role in the sentence. Specifically, it utilizes a dependency tree to calculate the importance of each word, i.e. the depth in the tree. A word far from the root is considered less important syntactically and is more likely to be removed (by means of dropout and blanking) or replaced during the augmentation process. Instead of using a syntax tree to assess word importance, Selective Text Augmentation (STA) [224] focuses more on the statistical and semantic features of words to describe the role of words. It proposes two perspectives to view words in text classification tasks, namely, statistical correlation, which means how frequently a word co-occurs with a certain category, and semantic similarity, which measures the semantic overlap of the given words and the category label. Using the metrics as axes, STA divides the roles of words into four quadrants: gold, venture, bonus, and trivial. Words with different roles are treated differently when imposing augmentation operations on the sentence. For example, gold words will be kept in selective deletion and synonyms of gold words will be added to the sentences in selective insertion. A summary of how these methods select target words is presented in Table IV.
TABLE IV: Word selection strategy of text data selection methods.
Name Ref Augmentation Strategy
@@
EDA [11]
- [43]
TEXTFOOLER [223]
LetterTran [225]
AEDA [46]
INSNET [226]
LeCA [227]
GBS [228]
- [47]
STA [224]
Subtree Swapping [156]
Substructure Substitution [157]
TreeMix [158]
SSMix [159]

### v4

@@ VI-A Vanilla Generation
A straightforward way to produce augmented data from the entire dataset is to learn the data distribution of existing datasets with neural networks and generate new samples that fall in that distribution. Some typical solutions are autoencoders (AE), generative adversarial networks (GAN), large language models (LLMs), and generative diffusion models.
@@ VI-A1 Autoencoder
An autoencoder learns from the dataset and extracts the essential structure, which can be used to generate a more robust dataset. It is widely used for tabular data augmentation. [168] tries out four types of autoencoder, concluding that all of these autoencoders can generate data sets that produce better task performance, while the variational autoencoder (VAE) provides the most robust result. Some other works [169, 170] also use VAE as their data generation approach. SDAT [169] adds extra noise to the latent space of its VAE to obtain augmented samples. These samples, along with the original ones, are then trained under a semi-supervised framework. [170] combines VAE with active learning. The samples are iteratively selected by active learning, labeled by an oracle, reconstructed by the VAE, and discriminated and classified. The use of AEs for feature space augmentation on image, text, and time-series data is discussed in [145], [16], and [171]. However, few follow-ups use such methods, potentially due to the inferior capacity compared with more advanced generation models. For instance, VAEs fail on complex datasets primarily because the limited approximation capacity of the encoder results in a poor, often overly simplistic, posterior distribution [172]. When applied to graph data, VAE models also require a massive graph-matching process [173].
@@ VI-A2 Generative Adversarial Networks
GANs enable generating new images on a large scale, leading to a new path of data augmentation.
For image data, GANs have now been a common data augmentation technique, especially in data-scarce scenarios like the medical field [174, 175]. Apart from simply enlarging the dataset, GAN can be used in other ways, such as removing background noise and cleaning the dataset [176, 177, 178]. It can also generate different views or poses of the object based on the original image [179], and thus add more variation to the dataset. Despite the huge success of GAN’s application for image data augmentation, a benchmark on biomedical image augmentation indicates that GAN-based methods do not generally outperform conventional data augmentation, very likely due to their limited generalizability [180].
Tabular data augmentation is another case where GANs play an important role. Table-GAN [181] mainly considers the privacy issue, and experiments show that the augmented table given by GAN is compatible with downstream machine learning models and produces similar performance. CTGAN [182] focuses on the class imbalance problem and proposes a conditional generator that accounts for the imbalance in categorical columns. [183] shares a similar goal and method with CTGAN [182] but employs a more sophisticated loss function that better incorporates the discriminator and the auxiliary classifier. ITS-GAN [184] proposes a framework that aims to preserve functional dependencies in tables. It models functional dependencies with autoencoders and includes a term to characterize the set of functional dependency constraints in the generator loss function. Apart from application to general tabular data, [185] extends the use of GAN to transcriptomics data, where each row of the table represents a sample, each column represents a gene, and the table’s values record the genes’ expression levels.
GANs are also widely used to generate virtual time-series data. [186] uses an RGAN and an RCGAN to produce realistic medical time-series data. TimeGAN [187] combines supervised and unsupervised losses to train its autoencoding components and adversarial components. [188] propose WaveGAN and SpecGAN to generate time-series data. WaveGAN is used on the waveform, while SpecGAN is used on the spectrogram. SpecGAN demonstrates how GAN is applied to generate time-series data in the frequency domain.
Although it is possible to generate graph data with GANs, [189] points out that GANs on graphs easily fall into mode collapse, where the generators tend to output non-unique samples, limiting GANs’ application on graph datasets.
@@ VI-A3 Large Language Models
Recent advances in large language models have established a new paradigm of data augmentation, especially benefiting the NLP field.
LAMBADA [190] is among the first attempts to use a state-of-the-art language generator to synthesize new text data for data augmentation. It generates a large dataset using a set of labels. Similarly, [191] uses the language model to generate a tree of tokens as candidates and constructs a sentence from them, and repeatedly refines the new sentences to build the augmented dataset. Data Boost [192] proposes a reinforcement learning framework to guide the generation process of GPT-2 to generate high-quality and high-similarity data. The rise of GPT-3 and later LLMs is now providing more possibilities for text generative augmentation, such as [193] and [194].
The development of multi-modal LLMs also facilitates image data augmentation. DIAGen [195] uses the GPT-4 model to provide image descriptions with additional meaningful context and uses the enriched textual descriptions to generate diverse image datasets. [196] designs a variety of textual prompts and utilizes the DALL-E model to produce an exclusively LLM-generated agricultural dataset, avoiding traditional fieldwork with imaging sensors.
LLMs can also be applied to tabular data augmentation by treating the table as text content. CLLM [197] uses the background description, in-context examples, and instruction guidelines to construct prompts and use the LLM to generate a large synthetic dataset, followed by a curator that evaluates and selects samples by their learning dynamics. Pred-LLM [198] first converts the table rows to text descriptions to fine-tune an LLM. Then it masks some features of sampled rows and uses the LLM to recover these rows. Finally, it uses the LLM to label the generated data.
Time-LLM [199] demonstrates the predictive ability of Transformer-based LLMs for time-series data, suggesting another possibility of LLMs as a data augmentation technique. However, beyond simply using LLMs to predict and expand sequences to future timestamps, the use of LLMs for general time-series data augmentation is still underexplored. Some exceptions are in the audio, voice, and speech fields, where the language information can be utilized by LLMs, enabling more possible manipulations, such as ArzEn-LLM [200], which employs the code-switched augmentations for multilingual speech recognition performance.
@@ VI-A4 Generative Diffusion Models
Diffusion models have recently emerged as a new solution to generative AI. Generating realistic and diverse samples by learning the underlying data distributions, they also show promising capabilities as a tool for data augmentation.
The diffusion model is widely applied to image data. GIF [201] leverages Stable Diffusion to optimize the latent features of the seed data in the semantically meaningful space of the prior model and create informative and realistic new data. SIGD [202] employs diffusion models to generate augmented images with good image diversity. It takes image labels and captions as guidance to maintain semantic consistency. [203] prompts the diffusion models to “generate a clean background” to avoid additional objects in the background and designs a background mask erosion mechanism to avoid extending objects to the background.
For text data, DiffusionCLS [204] leverages a diffusion LM to capture in-domain knowledge and generate pseudo samples by reconstructing from the corrupted dataset to balance between consistency and diversity. DiffLM [205] leverages diffusion models to reserve more information of the dataset’s original distribution and format structure in the learned latent distribution and decouples the learning of target distribution knowledge from the LLM’s generative objectives via a plug-and-play latent feature injection module.
The diffusion model also fills the gap of tabular data augmentation. [206] leverages diffusion models to first learn an unconditional generative model and introduces lightweight controllers to guide the unconditional generative model in generating synthetic data that satisfies different conditions. TabDiff [207] develops a joint continuous-time diffusion process for numerical and categorical data and proposes feature-wise learnable diffusion processes to counter the high disparity of different feature distributions. Forest-Diffusion [208] introduces a novel approach for generating and imputing mixed-type tabular data utilizing score-based diffusion and conditional flow matching.
Compared with other modalities, using the diffusion model for graph data and time-series data augmentation is relatively underexplored. To implement diffusion models on graphs, some properties of graphs must be properly treated. For example, diffusion models often rely on a continuous Gaussian noise process, but the graph structures are discrete [209]. In response, DiGress [210] designs a discrete denoising diffusion model for graph generation and achieves state-of-the-art performance on molecular and non-molecular datasets. Future advancement of generative diffusion models for graph data augmentation will especially benefit studies on molecules, proteins, and materials, whose data are often in the form of graph instances. For tabular data, [211] proposes ASE-DDPM for unbalanced time series classification. [212] employs the diffusion model for time-series data but only uses it to generate the first moment data. Despite these attempts, literature focusing on the design and application of diffusion-based time-series data augmentation is still scarce. Given the remarkable capability of diffusion models and their success in other data modalities, it is promising to further delve into this field.
@@ VI-B1 Introducing Exogenous Datasets
Tabular data populates the web, and by retrieving consistent and related tables, it is possible to generate an augmented table that incorporates extended information. A key consideration of table retrieval augmentation is to design appropriate ways to search for ideal tables. Santos [215] leverages external knowledge bases, YAGO 4, and focuses on designing a union search technique to match desired target tables. HYTREL [216] predicts table similarities by building hypergraphs on top of the tables and uses a hypergraph encoder to calculate the encoding. LEKA [217] uses an LLM to extract the table’s descriptive information, retrieve it in the Kaggle database, and harmonize tables from different sources.
@@ VI-B2 Leveraging Expert Knowledge
Computer Graphic Modeling is to leverage expert knowledge on how objects in the existing dataset look and the power of computer software to generate new image samples [218, 219]. These datasets have several advantages over ordinary image datasets. First, it requires little human effort to collect data. Second, by applying different aspects, poses, and lighting conditions, computer graphics modeling can generate images with high diversity. Third, these images are all fully labeled with accurate ground truth.
Relational Structure Construction can enhance the potential of tabular data, which is limited by its simple nature compared with data from other modalities. Due to the lack of spatial or relational relationships, many popular augmentation approaches and network configurations are not suitable for tabular datasets. A solution is to construct and introduce a relational structure based on the tabular dataset. GOGGLE [220] proposes a solution to this problem. It learns the relational structure and corresponding functional relationships from tabular datasets and uses this basis to generate new samples. In other words, it learns a graph from the features in the table and obtains augmented data from that graph. PET [221] depicts multiple tabular data instances as a hypergraph and uses the labels to help construct hyperedges. These methods are based on external assumptions and introduce extra structural information beyond the tabular dataset.
@@ VII Discussions
Up to now, this survey focuses on wherefrom the information in the augmented samples is derived. And there are many other perspectives to discuss data augmentation and understand how it works.
@@ VII-A Augmentation Evaluation Metrics
Most research on data augmentation uses the models’ performance on downstream tasks to evaluate the effect of data augmentation [14]. However, this criterion depends on many extraneous factors and may fail to reflect the influence of data augmentation methods on the data. So, there is a need for a set of evaluation metrics that can assess specific considerations of data augmentation. Two of the most critical metrics are affinity and diversity [224, 225, 226].
Affinity is also referred to as consistency, correctness, faithfulness, quality, similarity, or validity. It reflects if the augmented sample is semantically consistent with the original sample and preserves the original label. Model-centric evaluation [225] defines affinity as the ratio of model $m$’s performance on the augmented validation dataset $D^{\prime}_{val}$ and that on the original validation dataset $D_{val}$, given by $\mathit{affinity}=\text{ACC}(m,D^{\prime}_{val})/\text{ACC}(m,D_{val})$, where $\text{ACC}(m,D)$ stands for model $m$’s performance on the evaluation dataset $D$. Data-centric evaluation, on the other hand, defines affinity as how much the augmented dataset deviates from the original dataset in the feature space. For example, [224] proposes Fréchet BERT Distance (FBD) to evaluate the deviation of two textual datasets, defined as: $\mathit{FBD}=\sqrt{\|m_{1}-m_{2}\|^{2}_{2}+Tr(C_{1}+C_{2}-2(C_{1}C_{2})^{1/2}}$, where $m$ is the mean vector and $C$ is the covariance matrix.
Diversity reflects the breadth of patterns in the dataset, and exposing the model to a diverse training dataset can improve its robustness. [227] considers a dataset as an ellipsoid and defines its diversity as the isocontour radius of the ellipsoid: $\mathit{diversity}=(r_{1}{\cdot}r_{2}{\cdot}...{\cdot}r_{H})^{\frac{1}{H}}=\sqrt[H]{\prod^{H}_{i}\sigma_{i}}$, where $H$ is the dimension of the feature vectors and $\sigma_{i}$ is the standard deviation along the $i$-th feature. [228] proposes to evaluate dataset diversity by calculating the dispersion: $\mathit{disersion}=\sum_{x_{i},x_{j}}d(x_{i},x_{j})$, where $d(x_{i},x_{j})=1-\frac{f(x_{i}){\cdot}f(x_{j})}{\|f(x_{i})\|\|f(x_{j})\|}$.
Some empirical studies investigate how affinity and diversity influence the effectiveness of data augmentation, and their real-world experimental results demonstrate that model performance takes the most advantage of data augmentation that achieves both high affinity and diversity [225, 226, 228].
@@ VII-B1 On Selecting Pixels
For pixel erasing, selecting target pixels is a necessity before erasing them. Image-aware Random Erasing (IRE) [11] randomly selects an area in the image. Object-aware random erasing (ORE) [11] restricts random erasing to the object’s bounding box if its location is known. [229] proposed Region-aware Random Erasing (RRE), which introduces ORE with upper bounds for width and height in the bounding boxes, and then randomly erases the background. The same consideration also applies to image cropping (IC) and mixup, and some importance-based strategies are used to preserve important contents in the images. For IC, [73] uses entropy as the importance score to select the most important regions, and [74] uses eye-tracking technique to locate important content. For mixup, Co-mixup [230] and Puzzle Mix [231] both use saliency to decide the mixing strength of different regions. A summary of how these methods select target pixels is presented in Table IV.
TABLE IV: Region selection strategy for pixel erasing, image cropping, and mixing methods.
@@
Cutout [10], Image-aware RE [11], Region-aware RE [229], Hide-and-seek [38], Grid-Mask [39] Object-aware RE [11], Region-aware RE [229] Entropy-bases IC [73], Gaze-based IC  [74], Co-mixup [230], Puzzle mix [231]
@@ VII-B2 On Selecting Words
Most value-based text transformation methods randomly select the word to be replaced, added, and deleted, but some methods consider the role a word plays in the sentence when deciding whether to impose perturbation.
For example, TEXTFOOLER [232] ranks words by their importance and replaces the important words with their synonyms. Syntax-aware data augmentation [48] improves upon older methods by modifying words based on their grammatical role in a sentence. It uses a sentence’s structure (a dependency tree) to determine a word’s importance; words less central to the sentence’s core meaning are more likely to be changed or removed. Selective Text Augmentation (STA) [233] ignores grammatical structure and instead classifies words based on their statistical and semantic relationship to a category. It then groups words into four roles (like ”gold” for important words). During augmentation, it treats words differently based on their role; for example, it will keep ”gold” words but add synonyms for them. A summary of how these methods select target words is presented in Table V.
TABLE V: Word selection strategy of text data selection methods.
@@
EDA [12], [42], AEDA [46], [47] TEXTFOOLER [232], LeCA [234], GBS [235], STA [233], Substructure Substitution [152], Subtree Swapping [151], TreeMix [153] LetterTran [236], INSNET [237], SSMix [154]